OpenStudy (anonymous):

sketch the graph of y=4x^2-4x-3

OpenStudy (jack1):

hey, so same as last time, it's called a "quadratic equation" and takes the form: \[y = ax ^{2} + bx + c\]

OpenStudy (jack1):

so the key points you want for your sketch ar: y intercept (when x=0) x intercepts (when y=0) turning point (value of "c")

OpenStudy (jack1):

so if you factorise it, you can get your x intercepts: y = ( 2x + 1) ( 2x - 3 )

OpenStudy (jack1):

therefore your x intercepts occur when y = 0, so either (2x + 1) = 0 or (2x - 3) = 0 so x = -1/2 or x = 3/2 those are your intercepts (where your graph cuts the x axis)

OpenStudy (anonymous):

;)

OpenStudy (jack1):

@ nuh, thats completely incorrect

OpenStudy (anonymous):

?no it's right..and the other x intercept is 1

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i thought it 4x^3

OpenStudy (anonymous):

srry

OpenStudy (anonymous):

x intercepts are -0.5 and 1.5

OpenStudy (anonymous):

hey whats the quadratic equation?

OpenStudy (anonymous):

srry stupid mistake b4..:) read the equation incorrectely lol

OpenStudy (anonymous):